# Autonomous Navigation Of Quadrupeds Using Coverage Path Planning

### v1

@@ Abstract
This paper proposes a novel method of coverage path planning for the purpose of scanning an unstructured environment autonomously. The method uses the morphological skeleton of the prior 2D navigation map via SLAM to generate a sequence of points of interest (POIs). This sequence is then ordered to create an optimal path given the robot’s current position. To control the high-level operation, a finite state machine is used to switch between two modes: navigating towards a POI using Nav2, and scanning the local surrounding. We validate the method in a leveled indoor obstacle-free non-convex environment on time efficiency and reachability over five trials. The map reader and the path planner can quickly process maps of width and height ranging between [196,225] pixels and [185,231] pixels in 2.52 ms/pixel and 1.7 ms/pixel, respectively, where their computation time increases with 22.0 ns/pixel and 8.17 $\boldsymbol{\mu}$s/pixel, respectively. The robot managed to reach 86.5% of all waypoints over all five runs. The proposed method suffers from drift occurring in the 2D navigation map.
@@
Robotics, Quadruped, Autonomous Navigation, Planning, Inaccessible Environment, Unstructured Environment, Coverage Path Planning, LiDAR, SLAM, Morphological Skeletonization, Shortest Path, Dijkstra Algorithm, Graph Theory, Finite State Machine, Dead Reckoning, ROS 2
@@ I Introduction
Due to advancements in technology and miniaturization, in the past decade surface (or ground) robots, such as wheeled and legged robots, have been increasingly adopted for diverse operations in harsh and inaccessible environments. One of the key challenges in such environments is that the infrastructure to support diverse operations does not readily exist. These environments include, for example, disaster response [1, 2, 3], mining operations [4, 5], space exploration [6, 7, 8, 9], surveillance in remote locations [10, 11], or hazardous industries like nuclear power plant maintenance [12, 13].
In such complex environments, legged robots are more versatile and robust compared to wheeled robots than other surface robots such as wheeled rovers, and can adaptively navigate uneven, rugged, or soft terrain. Legged robots can cover relatively larger spatial areas by choosing safe footholds within their range of motion and rapidly responding to adjust their kinematic configuration [14] to achieve their objectives. The number of legs in a legged robot determines their movement efficiency and ability to maintain stability [15]. Compared to bipedal humanoids, quadruped robots demonstrate a greater load capacity and improved stability due to their broader base of support. On the other hand, quadrupeds possess simpler structures and control mechanisms than hexapodal and octopodal robots [16, 17]. For this reason, quadruped robots are ideal for tasks involving safe navigation of complex 3D environments for (sub-)surface exploration.
@@
There are several quadrupedal robots that are already commercially available in the market. We compare three notable examples, Boston Dynamics’ Spot, ANYbotics’ ANYmal, and Unitree Robotic’s Go2 Edu as seen in Figure 1, regarding attributes related to access of development, operation durability, and affordability, as seen in Table I. Both Spot and Anymal have garnered significant popularity and have made substantial contributions in research and engineering [18, 19] compared to that of Go2 Edu. However, their operational runtime is limited, and their cost is considerably higher. Go2 has three modes: Air, Pro, and Edu with a cost of 1600 USD, 2800 USD, and 12500 USD, respectively. However, only the Edu mode allows for software development which is necessary for custom implementations, including other necessary features. Furthermore, Go2 Edu has a dedicated Robotics Operating System 2 (ROS 2) integration that allows rapid development and testing. With a factor of two to three in running time, experiments can be done over a long session and the robot can provide sufficient battery capacity to power additional sensors. This paper focuses on the development of a software architecture for path planning and navigation specific to Go2 Edu.
@@ I-A Related Works
Autonomous navigation using quadrupedal robots is crucial for exploring complex environments; however, limited work has been done on 2D coverage path planning. The work in [20] presented a 2D coverage path planner for investigating unknown and unstructured environments while accounting for time-bounded and dynamics constraints, and traversability risk. Research in [21] considered the 2D coverage path planning as Travelling Salesman Problem (TSP) by applying integer linear programming, where environment perception was achieved using RGB-D camera inputs. In another recent work [22], a novel autonomous exploration method was developed using topological skeleton of the environment’s geometry via LiDAR and a finite state machine (FSM) that enables exploratory strategy. This was demonstrated on a quadrupedal robot in an unstructured environment. These relevant works prove that quadrupedal robots are suitable to cover the whole environment and indicate the feasibility of achieving 2D coverage path planning using the Unitree Robotics Go2 Edu robot.
Several studies have recently explored the development of software infrastructure for the Unitree Robotics Go2 Edu robot. In [23], a reinforcement learning method was developed to enable the Go2 Edu robot to navigate narrow pipes using visual inputs from a depth camera. The navigation process is relatively straightforward due to the grid-like structure of the pipes, where each pipe is aligned in a straight path with occasional protrusions serving as obstacles. The work in [24] developed a high-level path planning using a large language model namely OpenAI’s ChatGPT-4o which allows the interpretation of human verbal commands and translates them into a list of executable instructions. The system integrates a depth camera alongside a segmentation model to perceive the environment effectively. In addition, research in [25] developed a motion controller for the Go2 Edu robot to traverse complex and unstructured environments using proprioceptive sensing and collision estimation only. Nevertheless, the missing aspect is the state of these recent works is to enable the Go2 Edu robot to systematically explore every reachable corner of an unstructured and unpredictable environment. Such a capability is an immediate need and would be beneficial for inspecting large and complex environments where manual inspection is impractical and labor-intensive. To the best of our knowledge, these innovations and experiments have not been reported in state-of-the-art literature.
@@ I-B Contribution
    Using a prior 2D navigation map of the surroundings, the framework rapidly generates an efficient path which ensures coverage from a small to a large area.
@@
    We validate the developed system with on field experiments on navigational tasks using the Unitree Robotics Go2 Edu robot in an indoor environment.
@@
    We implemented an extended interface that enables a Unitree Robotics Go2 Edu to work together with the control framework using ROS 2.
@@ II-D Navigation
We use the Nav2 framework [30] to ensure the robot is able to plan and navigate towards a desired pose (position and orientation) in the environment, which is also completely compatible with the SLAM Toolbox discussed previously. Using the 2D navigation map provided by SLAM, the Nav2 framework can plan and navigate a path as a function of the 2D poses ($x$- and $y$-positions, and orientation about the $z-$ axis $\theta$) of the destination, the robot’s current location, and the robot’s kinematics constraints with respect to its surroundings.
The Nav2 framework includes an array of tools such as planners, recoveries and controllers. It is outside the scope of this paper to experiment different tools. Therefore, we mainly used the default configuration with minor adjustments that corresponds to the robot’s kinematics. It we can consider the robot to behave similar to a differential wheeled robot. This is due to the fact that the robot’s movement is controlled using 2D velocity $\dot{x},\dot{y},\dot{\theta}$ as input. With this in mind, we can simply use Nav2’s default planner NavFnPlanner.
@@ III-A Map Reader
To effectively cover the environment, the robot must follow a path that ensures it traverses every corner of the space. This path should align with the trajectory of the occupied areas e.g. corridors. This approach is generally effective under the assumption that the environment is confined within a bounded, occupied space, such as an indoor setting. In order to create the path, we process the 2D navigation map as an 8-bit array using the algorithm as seen in Algorithm 1. $\mathbf{M}$, $H$, $W$ denote the map, and the map’s height and width, respectively. The map is divided into cells using the map resolution $R$ in m/pixel. Each of these cells is an element in $\mathbf{M}$, $m_{ij}$ which only contains one of the three types of space: occupied, free, and $unknown$, which are represented by the integers 0, 255, and 128, respectively.
@@
2:Input: Hyperparameters $\sigma,\kappa,\mathbf{K}$
@@
4:for $i\in H$ do # Replace unknown cells as occupied
5:     for $j\in W$ do
@@
26:              $\mathcal{V}\leftarrow\mathcal{V}\cup\begin{bmatrix}i&j\end{bmatrix}^{T}$
@@
30:return $\mathcal{V}\in\mathbb{R}^{O\times 2}$
Algorithm 1 Process the 8-bit 2D navigation map $\mathbf{M}$ into a set of 2D waypoints $\mathcal{V}$ where each element is a waypoint $\boldsymbol{v}_{i}\in\mathbb{R}^{2}$.
Firstly, the majority of the unknown spaces in the map are located outside the boundaries of the occupied area e.g. beyond walls. Consequently, these spaces are therefore inaccessible and we can simply treat the unknown regions as occupied space. To establish a smooth connection between occupied cells affected by noise in the map, a Gaussian filter[31] with a standard deviation $\sigma\in\mathbb{N}$ is applied. Subsequently, the result is binarized using a threshold value $\kappa\in[0,255]$. Due to remaining noise and transparent obstacles such as glass, this means that there may be free space that is unreachable to the robot. To mitigate this issue, we assume that the path is always the largest 2D contour. Firstly, the contours are found using marching cubes algorithm [32]. Afterwards, we recreate the map by filling the largest 2D contour with 255. However, it is worth noting that by only filling the largest contour, it does not take obstacles inside the contour into account.
@@
To ensure that the robot remains a safe distance from the occupied space during navigation, the free space in the map is shrunk using a morphological filter called erosion [33]. Erosion uses a structuring element $\mathbf{K}$, also referred to as a kernel, that determines the width to be removed. For the sake of simplicity, we take a square matrix of ones as our structuring element: $\mathbf{K}=\boldsymbol{1}_{k}\boldsymbol{1}_{k}^{T}$, where $k$ determines the length of the vector and $k=1,2,3,\ldots$. The larger the $k$, the larger the safe distance. In addition, the safe distance reduces when $R$ reduces. Therefore, $k\propto R$. This also removes narrow passages that prove impassable for the robot.
@@
The remaining free space is reduced to a thin one-pixel-wide representation containing $O$ pixels, which corresponds to the topological skeleton of the map’s geometry. In a practical scenario, the time complexity of 2D skeletonization is mainly proportional to the total number of pixels in an image $P$ as it iterates until the object becomes one-pixel wide. Therefore it can be considered as $\mathcal{O}(P)$ [34]. This 2D skeleton map $\mathbf{M_{\mathbf{S}}}$ is then flattened into an unordered set that describe the $x$- and $y$-position of every 2D point $\mathcal{V}$, where their corresponding pixels satisfies $m_{\mathbf{S},ij}=255$. Given that the skeleton contains $O$ pixels, it follows that there are $O$ waypoints in $\mathcal{V}$. We refer every such 2D point defined by the skeleton as waypoints, denoted as $\boldsymbol{v}_{i}\in\mathbb{R}^{2}$ where $0\leq i\leq O$. The waypoints serves as points of interests for the robot to visit.
@@ III-B Path Planning
In order to scan the whole environment, the robot should perform the scanning procedure for each waypoint. Therefore, the objective is to determine a path that includes every waypoint the robot must visit while optimizing for time efficiency. Ultimately, this can be considered as a Travelling Salesman Problem (TSP) which we approach by formulating a fast and efficient path planning algorithm. It takes a graph $\mathcal{G}$ as input, and outputs a path $\mathcal{P}$. $\mathcal{P}$ is the ordered set of waypoints, in other words, a sequence of vertices $\mathbf{v}_{i}$ traversed by the robot. $\mathcal{G}$ is constructed from the unordered set of waypoints obtained from map reader $\mathcal{V}$ as vertices. These vertices are then connected to other vertices if they are neighbors around the map’s resolution $R$. This should resemble to the skeleton representation of the map, where the connections define the edges $\mathcal{E}$ of $\mathcal{G}$. The pseudocode can be seen in Algorithm 2.
@@
3:$\mathcal{V}_{leaf,total}\leftarrow\emptyset$
4:for $\boldsymbol{v}_{i}\in\mathcal{V}$ do # Gather all leaf vertices in a set
@@
6:         $\mathcal{V}_{leaf,total}\leftarrow\mathcal{V}_{leaf,total}\cup\boldsymbol{v}_{i}$
@@
9:$\mathbf{v}_{leaf,start}\leftarrow\text{FindNearestLeaf}(\boldsymbol{x}_{0},\mathcal{V}_{l,total})$
@@
11:$\mathcal{V}_{leaf,visited}\leftarrow\emptyset$
12:while $|\mathcal{V}_{leaf,visited}|<|\mathcal{V}_{leaf,total}|$ do
13:     $\mathcal{V}_{leaf,visited}\leftarrow\mathcal{V}_{leaf,visited}\cup\boldsymbol{v}_{leaf,start}$
14:     $\boldsymbol{v}_{leaf,target}\leftarrow\text{FindNearestLeaf}(\boldsymbol{v}_{leaf,sstart},\mathcal{V}_{leaf,total}\setminus\boldsymbol{v}_{leaf,visited})$
15:     $\mathcal{P}^{\prime}\leftarrow\text{FindShortestPath}(\boldsymbol{v}_{leaf,start},\boldsymbol{v}_{leaf,target})$
16:     for $\boldsymbol{v}^{\prime}_{i}$ in $\mathcal{P}^{\prime}$ do
@@
18:              $\mathcal{P}\leftarrow\mathcal{P}\cup\boldsymbol{v}^{\prime}_{i}$
@@
21:     $\boldsymbol{v}_{leaf,start}\leftarrow\boldsymbol{v}_{leaf,target}$
@@
24:for $k\leftarrow 0$ to $\left\lfloor{\frac{|\mathcal{P}|-1}{D/R}}\right\rfloor$ do # Downsample the path
@@
27:return $\mathcal{P}^{*}\in\mathbb{R}^{2}$
Algorithm 2 A fast and efficient approach to planning a path using the unordered set of waypoints obtained from map reader $\mathcal{V}$, map resolution $R$, waypoint resolution $D$, and the robot’s starting 2D position $\boldsymbol{x}_{0}$.
First of all, there are dead-ends found in the skeleton representation of the map. For the robot to cover the whole environment, we can utilize these dead-ends. Each dead-end is identified as a leaf vertex $\boldsymbol{v}_{leaf}\in\mathcal{V}_{leaf}$ and are defined with degree 1. So to fully cover the map, the robot should visit every leaf vertex in $\mathcal{V}_{leaf}$ sequentially. Visiting from one leaf vertex to another also enables the robot to visit unvisited vertices of degree 2 or higher.
After identifying all leaf vertices, the leaf vertex closest to the robot’s current 2D position $\boldsymbol{x}_{0}$ is selected as the initial source leaf vertex $\boldsymbol{v}_{leaf,start}$ using Algorithm 3. For every source leaf vertex $\boldsymbol{v}_{leaf,start}$, we find the next nearest unvisited leaf vertex as the target $\boldsymbol{v}_{leaf,target}$ using Algorithm 3. We find for the shortest path $\mathcal{P}^{\prime}$ between the the source $\boldsymbol{v}_{leaf,source}$ and the target $\boldsymbol{v}_{leaf,target}$ in the graph. To achieve this, we use Dijkstra’s algorithm, which has a time complexity of $\mathcal{O}(|\mathcal{E}|+|\mathcal{V}|\log|\mathcal{V}|)$, where $|\mathcal{V}|$ and $|\mathcal{E}|$ denote the total number of nodes and the total number of edges in the graph, respectively [35]. Once $\mathcal{P}^{\prime}$ is found, each vertex $\boldsymbol{v}^{\prime}_{i}\in\mathcal{P}^{\prime}$ is appended into $\mathcal{P}$. Unless, if some vertex $\boldsymbol{v}^{\prime}_{i}$ appears again but from another $\mathcal{P}^{\prime}$, it is not worth wasting time scanning at this position again. Thus, instead it will be ignored. Once iterated over all $\boldsymbol{v}^{\prime}_{i}$ in $\mathcal{P}^{\prime}$, $\boldsymbol{v}_{leaf,start}$ is inserted into a set of visited leaf vertices $\mathcal{V}_{leaf,visited}$ and $\boldsymbol{v}_{l,t}$ becomes the next source $\boldsymbol{v}_{leaf,start}$. Not only is the shortest path useful for appending $\boldsymbol{v}_{i}$ into $\mathcal{P}$ but also for ensuring a feasible and unobstructed trajectory through occupied spaces.
@@
2:Input: Source 2D position $\boldsymbol{x}_{0}\in\mathbb{R}^{2}$, a selected set of leaf nodes $\mathcal{V}^{\prime}_{leaf}$
3:$\boldsymbol{v}_{leaf,target}\leftarrow\emptyset$
@@
5:for $\mathbf{v}_{leaf,i}\in\mathcal{V}^{\prime}_{leaf}$ do
6:     if $|\boldsymbol{v}_{leaf,i}-\boldsymbol{x}_{0}|<d$ then
7:         $\boldsymbol{v}_{leaf,target}\leftarrow\boldsymbol{v}_{leaf,i}$
8:         $d\leftarrow|\boldsymbol{v}_{leaf,i}-\boldsymbol{x}_{0}|$
@@
11:return $\boldsymbol{v}_{leaf,target}$

### v2

@@ Abstract
[…]
@@
Quadruped, Autonomous Navigation, Unstructured Environment, Coverage Path Planning, ROS 2
@@ I Introduction
Due to advancements in technology and miniaturization, in the past decade surface (or ground) robots, such as wheeled and legged robots, have been increasingly adopted for diverse operations in harsh and unstructured environments. One of the key challenges in such environments is that the infrastructure to support diverse operations does not readily exist. These environments include, for example, disaster response [1, 2, 3], mining operations [4, 5], space exploration [6, 7, 8, 9], surveillance in remote locations [10, 11], or hazardous industries like nuclear power plant maintenance [12, 13].
In such complex environments, legged robots are more versatile and robust compared to wheeled robots than other surface robots such as wheeled rovers, and can adaptively navigate uneven, rugged, or soft terrain. Legged robots can cover relatively larger spatial areas by choosing safe footholds within their range of motion and rapidly responding to adjust their kinematic configuration [14] to achieve their objectives. The number of legs in a legged robot determines its movement efficiency and ability to maintain stability [15]. Compared to bipedal humanoids, quadruped robots demonstrate a greater load capacity and improved stability due to their broader base of support. On the other hand, quadrupeds possess simpler structures and control mechanisms than hexapodal and octopodal robots [16, 17]. For this reason, quadruped robots are ideal for tasks involving safe navigation of complex 3D environments for (sub-)surface exploration.
@@
There are several quadrupedal robots that are already commercially available on the market. We compare three notable examples, Boston Dynamics’ Spot, ANYbotics’ ANYmal, and Unitree Robotic’s Go2 Edu as seen in Figure 1, regarding attributes related to access of development, operation durability, and affordability, as seen in Table I. Both Spot and Anymal have garnered significant popularity and have made substantial contributions in research and engineering [18, 19] compared to that of Go2 Edu. However, their operational runtime is limited, and their cost is considerably higher. Go2 has three modes: Air, Pro, and Edu with a cost of 1600 USD, 2800 USD, and 12500 USD, respectively. However, only the Edu mode allows for software development, which is necessary for custom implementations, including other necessary features. Furthermore, Go2 Edu has a dedicated Robotics Operating System 2 (ROS 2) integration that allows rapid development and testing. With a factor of two to three in running time, experiments can be done over a long session and the robot can provide sufficient battery capacity to power additional sensors. This paper focuses on the development of a software architecture for path planning and navigation specific to Go2 Edu.
The goal of this research is to enable quadruped robots to map the terrain using coverage path planning. For quadrupeds to do this, they require sensors, where the most common are cameras [20, 21, 22] and LiDAR [23, 24]. To be able to scan with the sensors, the robot needs to move to various POIs, which requires coverage path planning and navigation methods. These will be examined in the related works below.
@@ I-A Related Works
Several studies have recently explored the development of software infrastructure for the Unitree Robotics Go2 Edu robot. In [25], a reinforcement learning method was developed to enable the Go2 Edu robot to navigate narrow pipes using visual inputs from a depth camera. The navigation process is relatively straightforward due to the grid-like structure of the pipes, where each pipe is aligned in a straight path with occasional protrusions serving as obstacles. The work in [26] developed a high-level path planning using a large language model, namely OpenAI’s ChatGPT-4o, which allows the interpretation of human verbal commands and translates them into a list of executable instructions. The system integrates a depth camera alongside a segmentation model to perceive the environment effectively. In addition, research in [27] developed a motion controller for the Go2 Edu robot to traverse complex and unstructured environments using proprioceptive sensing and collision estimation only.
Autonomous navigation using quadrupedal robots is crucial for exploring complex environments; however, limited work has been done on 2D coverage path planning. Research in [28] achieved coverage path planning for loco-manipulation through an integrated end-to-end pipeline combining perception, optimization, and whole-body motion planning with RGB-D camera inputs. Another work in [23] presented a 2D coverage path planner for investigating unknown and unstructured environments while accounting for time-bounded and dynamics constraints, and traversability risk. The advantage of these approaches is that they guarantee timely execution when the mission is time-bound, and they find a good optimal tradeoff between maximum area coverage and path traveled. The drawback of that approach is that it is focused on time constraints, therefore, it does not work when the task has a variable execution time and requires time to be defined in advance. In contrast, our approach enables planning for variable times.
Some navigation methods already enable planning for variable times. In the recent work [24], a novel autonomous exploration method was developed using a topological skeleton of the environment’s geometry via LiDAR and a finite state machine (FSM) that enables exploratory strategy. This was demonstrated on a quadrupedal robot in an unstructured environment. The advantage of these approaches is very adaptable in an unknown environment due to the use of a state machine that continuously checks for undiscovered areas within the map. The main difference between this work and our approach is that their method for obtaining topological skeletons uses wave propagation and Voronoi diagrams, while our method uses a morphological technique by treating the environment as an image. The impact of this is that the morphological technique simplifies the mapping and is thus computationally more efficient.
Furthermore, all the above methods on coverage path planning have specific map generation algorithms that are an integral part of the whole approach. Therefore, they are less modular, and it is not easy to replace them with developing state-of-the-art mapping algorithms. In contrast, our approach can work with different types of mapping algorithms, thus making it more modular.
@@ I-B Contribution
    Using a prior 2D navigation map of the surroundings, the framework rapidly generates an efficient path based on the morphological skeleton of the map, which ensures coverage from a small to a large area.
@@
    We validate the developed system with on-field experiments on navigational tasks using the Unitree Robotics Go2 Edu robot in an indoor environment.
@@
    We designed an extended interface that enables a Unitree Robotics Go2 Edu to work together with the control framework using ROS 2.
@@ II-D Navigation
We use the Nav2 framework [33] to ensure the robot is able to plan and navigate towards a desired pose (position and orientation) in the environment, which is also completely compatible with the SLAM Toolbox discussed previously. Using the 2D navigation map provided by SLAM, the Nav2 framework can plan and navigate a path as a function of the destination’s pose, the robot’s current pose, and the robot’s kinematics constraints with respect to its surroundings. A 2D pose is defined as $\boldsymbol{x}:=\begin{bmatrix}x&y&\psi\end{bmatrix}^{T}$, where $x$ and $y$ define the longitudinal and lateral displacements, and $\psi$ as the orientation of the robot about the $z$-axis (vertical displacement).
The Nav2 framework includes an array of tools such as planners, recoveries and controllers. It is outside the scope of this paper to experiment with different tools. Therefore, we mainly used the default configuration with minor adjustments that corresponds to the robot’s kinematics. Due to the fact that the robot’s movement can be controlled using 2D velocities $\dot{x},\dot{y},\dot{\theta}$ as inputs, we can consider the robot to behave similarly to that of a differential wheeled robot. With this in mind, we can simply use Nav2’s default planner NavFnPlanner.
@@ III-A Map Reader
To effectively cover the environment, the robot must follow a path that ensures it traverses every corner of the space. This path should align with the trajectory of the occupied areas, e.g. corridors. This approach is generally effective under the assumption that the environment is confined within a bounded, occupied space, such as an indoor setting. In order to create the path, we process the 2D navigation map as an 8-bit array using the algorithm as seen in Algorithm 1. $\mathbf{M}$, $H$, $W$ denote the map, and the map’s height and width, respectively. The map is divided into cells using the map resolution $R$ in m/pixel. Each of these cells is an element in $\mathbf{M}$, $m_{ij}$, which only contains one of the three types of space: occupied, free, and $unknown$, which are represented by the integers 0, 255, and 128, respectively.
Algorithm 1 Process the 8-bit 2D navigation map $\mathbf{M}$ into an unordered set of waypoints $\mathcal{V}$ where each element is a waypoint $\boldsymbol{v}\in\mathbb{R}^{2}$.
@@
2:Input: Hyperparameters $\sigma,\kappa,\mathbf{K},\boldsymbol{o}\in\mathbb{R}^{2},R$
@@
4:for $i\in\{1,\ldots,H\}$ do
5:     for $j\in\{1,\ldots,W\}$ do
@@
26:              $\mathcal{V}\leftarrow\mathcal{V}\cup\{R\begin{bmatrix}i&j\end{bmatrix}^{T}+\boldsymbol{o}\}$
@@
30:return $\mathcal{V}$
First, consider an indoor environment. In such case, most unknown cells inside the map are located outside the boundaries of the occupied area e.g. beyond walls. This is because SLAM initially represents the environment as a grid of unknown cells before any exploration occurs. Furthermore, unknown cells often persist between occupied and free regions as a result of occlusions encountered during raycasting-based measurements, such as those performed using LiDAR, as seen in Figure 6(a). Consequently, this unknown space cannot be reached by the robot and we can simply treat the unknown regions as part of the boundaries–the occupied space. This can be seen in Figure 6(b).
To establish a smooth connection between occupied cells affected by noise in the map, a Gaussian filter [34] is applied using a standard deviation parameter $\sigma\in\mathbb{N}$, as seen in Figure 6(c). Subsequently, in order to restore the binary representation of occupied and the free cells, the filtered result is binarized using a threshold parameter $\kappa\in[0,255]$. A cell with a value greater than $\kappa$ is classified as a free cell; otherwise, it is designated as an occupied cell.
Due to residual noise and the presence of transparent obstacles, such as windows, this means that there may be free cells that is unreachable to the robot. To mitigate this issue, we assume that the navigable space for the robot corresponds the largest 2D contour. First, the contours within the map are found using the marching cubes algorithm [35]. The map is then reconstructed by filling the largest 2D contour with a value of 255, as seen in Figure 6(d). It is worth noting that by only filling the largest contour, it does not take obstacles inside the contour into account.
@@
To ensure that the robot remains a safe distance from the occupied space during navigation, the free space in the map is shrunk using a morphological filter called erosion [36], as seen in Figure 6(e). Erosion uses a structuring element $\mathbf{K}$, also referred to as a kernel, that determines the width to be removed. For the sake of simplicity, we take a square matrix of ones as our structuring element: $\mathbf{K}=\boldsymbol{1}_{k}\boldsymbol{1}_{k}^{T}$, where $k$ determines the length of the vector and $k=1,2,3,\ldots$. The larger the $k$, the larger the safe distance. In addition, the safe distance reduces when $R$ reduces. Therefore, $k\propto R$. This also removes narrow passages that prove impassable for the robot.
@@
The remaining free space is reduced to a thin one-pixel-wide representation containing $O$ pixels, which corresponds to the topological skeleton of the map’s geometry. This can be seen in Figure 6(f). In a practical scenario, the time complexity of 2D skeletonization is mainly proportional to the total number of pixels in an image $P$ as it iterates until the object becomes one-pixel wide. Therefore, it can be considered as $\mathcal{O}(P)$ [37].
The 2D skeleton map $\mathbf{M_{\mathbf{S}}}$ is then flattened into an unordered set that describe the $x$- and $y$-position of every 2D point by transforming the pixel coordinates into real-world coordinates using the origin $\boldsymbol{o}$ and resolution $R$ of the 2D navigation map. Only those pixels for which $m_{\mathbf{S},ij}=255$ are considered in this transformation. Consider that the skeleton contains $O$ pixels, it follows that there are $O$ waypoints in $\mathcal{V}$. We refer to every such 2D point defined by the skeleton as waypoints, denoted as $\boldsymbol{v}_{i}\in\mathbb{R}^{2}$ where $0\leq i\leq O$. The waypoints serve as points of interests for the robot to visit.
@@ III-B Path Planning
In order to scan the whole environment, the robot should perform the scanning procedure for each waypoint. Therefore, the objective is to determine a path that includes every waypoint the robot must visit while optimizing for time efficiency. Ultimately, this can be considered as a Travelling Salesman Problem (TSP) which we approach by formulating a fast and efficient path planning algorithm. It takes a connected acyclic graph $\mathcal{G}$ as input, and outputs a path $\mathcal{P}$, which is the ordered set of waypoints. In other words, $\mathcal{P}$ is a sequence of vertices $\boldsymbol{v}_{i}$ traversed by the robot.
In order to construct the graph $\mathcal{G}$, we treat the unordered set of waypoints $\mathcal{V}$, obtained from the map reader, as the set of vertices. Each vertex in $\mathcal{V}$ uniquely corresponds to the coordinate vector of a waypoint, and can, for the sake of simplicity, be denoted as $\boldsymbol{v}_{i}\in\mathbb{R}^{2},\;\forall i=1\ldots,O$. Each vertex is then connected to other vertices if they are neighbors around the map’s resolution $R$. This should resemble to the skeleton representation of the map, where the connections define the edges $\mathcal{E}$ of the graph $\mathcal{G}$.
@@
3:$\mathcal{V}_{\text{leaf,total}}\leftarrow\emptyset$
4:for $\boldsymbol{v}_{i}\in\mathcal{V}$ do
@@
6:         $\mathcal{V}_{\text{leaf,total}}\leftarrow\mathcal{V}_{\text{leaf,total}}\cup\{\boldsymbol{v}_{i}\}$
@@
9:$\mathcal{V}_{\text{leaf,visited}}\leftarrow\emptyset$
@@
11:$\boldsymbol{v}_{\text{leaf,start}}\leftarrow\text{FindNearestLeaf}(\boldsymbol{x}_{0},\mathcal{V}_{\text{leaf,total}})$
12:while $|\mathcal{V}_{\text{leaf,visited}}|<|\mathcal{V}_{\text{leaf,total}}|$ do
13:     $\mathcal{V}_{\text{leaf,visited}}\leftarrow\mathcal{V}_{\text{leaf,visited}}\cup\{\boldsymbol{v}_{\text{leaf,start}}\}$
14:     $\boldsymbol{v}_{\text{leaf,target}}\leftarrow\text{FindNearestLeaf}(\boldsymbol{v}_{\text{leaf,start}},\mathcal{V}_{\text{leaf,total}}\setminus\mathcal{V}_{\text{leaf,visited}})$
15:     $\mathcal{P}^{\prime}\leftarrow\text{FindShortestPath}(\boldsymbol{v}_{\text{leaf,start}},\boldsymbol{v}_{\text{leaf,target}}\mid\mathcal{G})$
16:     for $\boldsymbol{v}^{\prime}_{i}\in\mathcal{P}^{\prime}$ do
@@
18:              $\mathcal{P}\leftarrow\mathcal{P}\cup\{\boldsymbol{v}^{\prime}_{i}\}$
@@
21:     $\boldsymbol{v}_{\text{leaf,start}}\leftarrow\boldsymbol{v}_{\text{leaf,target}}$
@@
24:for $k\leftarrow 0$ to $\left\lfloor{\frac{|\mathcal{P}|-1}{D/R}}\right\rfloor$ do
@@
27:return $\mathcal{P}^{*}$
Algorithm 2 A fast and efficient approach to planning a path $\mathcal{P}^{*}$, with $L$ number of vertices, using the unordered set of waypoints obtained from map reader $\mathcal{V}$, map resolution $R$, waypoint resolution $D$, and the robot’s starting 2D position $\boldsymbol{x}_{0}$.
Algorithm 2 outlines the procedure for obtaining an efficient path $\mathcal{P}^{*}$ for coverage exploration of the whole environment. First, there are dead-ends found in the skeleton representation of the map. For the robot to cover the whole environment, we can utilize these dead-ends. Each dead-end is identified as a leaf vertex $\boldsymbol{v}_{\text{leaf}}\in\mathcal{V}_{\text{leaf}}$ and is typically defined with degree 1. Assuming the graph $\mathcal{G}$ is connected and acyclic, complete coverage of the map can be achieved by ensuring that the robot visits every leaf vertex in $\mathcal{V}_{\text{leaf}}$ in sequence. Since all other vertices in the graph lie on the paths connecting the leaves. traversing to each leaf inherently requires passing through the intermediate vertices. As a consequence, visiting all leaf vertices implies that the entire graph has been traversed.
After identifying all leaf vertices, the leaf vertex closest to the robot’s current 2D position $\boldsymbol{x}_{0}$ is selected as the initial source leaf vertex $\boldsymbol{v}_{\text{leaf,start}}$ using Algorithm 3. For every source leaf vertex $\boldsymbol{v}_{\text{leaf,start}}$, we find the next nearest unvisited leaf vertex as the target $\boldsymbol{v}_{\text{leaf,target}}$ using FindNearestLeaf, as defined in Algorithm 3. We find the shortest path $\mathcal{P}^{\prime}$ between the source $\boldsymbol{v}_{\text{leaf,source}}$ and the target $\boldsymbol{v}_{\text{leaf,target}}$ given the graph $\mathcal{G}$. To achieve this, we use the method FindShortestPath which essentially utilizes Dijkstra’s algorithm. This algorithm has a time complexity of $\mathcal{O}(|\mathcal{E}|+|\mathcal{V}|\log|\mathcal{V}|)$, where $|\mathcal{V}|$ and $|\mathcal{E}|$ denote the total number of vertices and the total number of edges in the graph, respectively [38].
Once $\mathcal{P}^{\prime}$ is found, each vertex $\boldsymbol{v}^{\prime}_{i}\in\mathcal{P}^{\prime}$ is appended into $\mathcal{P}$. Unless some vertex $\boldsymbol{v}^{\prime}_{i}$ reappears as a path of a different path $\mathcal{P}^{\prime}$, revisiting and scanning at this position is unnecessary and should be avoided to conserve time and computational resources. Once iterated over all $\boldsymbol{v}^{\prime}_{i}$ in $\mathcal{P}^{\prime}$, $\boldsymbol{v}_{\text{leaf,start}}$ is inserted into a set of visited leaf vertices $\mathcal{V}_{\text{leaf,visited}}$ and $\boldsymbol{v}_{\text{leaf,target}}$ becomes the next source $\boldsymbol{v}_{\text{leaf,start}}$. Not only is the shortest path useful for appending $\boldsymbol{v}_{i}$ into $\mathcal{P}$ but also for ensuring a feasible and unobstructed trajectory through occupied spaces.
@@
2:Input: $\boldsymbol{x}_{0}\in\mathbb{R}^{2}$, $\mathcal{V}^{\prime}_{\text{leaf}}$
3:$\boldsymbol{v}_{\text{leaf,nearest}}\leftarrow\boldsymbol{0}$
@@
5:for $\boldsymbol{v}_{\text{leaf,i}}\in\mathcal{V}^{\prime}_{\text{leaf}}$ do
6:     if $|\boldsymbol{v}_{\text{leaf,i}}-\boldsymbol{x}_{0}|<d$ then
7:         $\boldsymbol{v}_{\text{leaf,nearest}}\leftarrow\boldsymbol{v}_{\text{leaf,i}}$
8:         $d\leftarrow|\boldsymbol{v}_{\text{leaf,i}}-\boldsymbol{x}_{0}|$
@@
11:return $\boldsymbol{v}_{\text{leaf,nearest}}$